Find all real solutions of the equation (2 – 6)? = 4. 21 = and 12 with i < 22
Answer #1

4 (ox - 6)2 = 4 → (2-6) = (432 =) 4-6=2 or x-6 = -2 = ne=8 or X =4 X;=8 & ng=4 csscanned with CamScanner